# Information Physics of Intelligence: Unifying Logical Depth and Entropy under Thermodynamic Constraints

### v3

@@ IV-C Entropy Production and Irreversibility
Theorem 11 (Minimum Entropy Production).
Answering query $q$ produces minimum environmental entropy:
$\Delta S_{\text{env}}\geq\frac{I(S;q)}{T}$ (89)
By the second law of thermodynamics, total entropy change:
$\Delta S_{\text{total}}=\Delta S_{\text{system}}+\Delta S_{\text{env}}\geq 0$ (90)
The system extracts information $I(S;q)$ from storage $S$. This reduces system entropy by $I(S;q)/\ln(2)$ nats. To maintain non-negative total entropy:
$\Delta S_{\text{env}}\geq\frac{I(S;q)}{T}$ (91)
V Optimal Hybrid Strategies
V-A Threshold Policies for Hybrid Strategies
Theorem LABEL:thm:duality implies that the marginal benefit of storing additional information diminishes as the knowledge base grows. For a query $q$ with access frequency $f_{q}$, the expected cost of computing on-demand is
$\mathbb{E}[\text{Compute Cost}]=f_{q}\cdot H_{\mathrm{derive}}(q\mid S_{O})\\ =\Theta\!\left(f_{q}\cdot\frac{H_{q}}{\log|\text{Atom}(S_{O})|}\right),$ (92)
while the amortized cost of storing its answer is
$\mathbb{E}[\text{Store Cost}]=H_{q}+\lambda\cdot\text{(storage overhead)},$ (93)
where $\lambda$ penalizes storage capacity usage.
A query should be stored if and only if
$f_{q}\cdot\frac{H_{q}}{\log|\text{Atom}(S_{O})|}>H_{q}\quad\Longleftrightarrow\quad f_{q}>\Theta(\log|\text{Atom}(S_{O})|).$ (94)
Thus, we obtain a scale-aware threshold policy:
Policy 1 (Scale-Aware Threshold).
Store query $q$ if its normalized frequency satisfies
$f_{q}\cdot\frac{Dd(q\mid S_{O})}{H_{q}}>\tau,$ (95)
where the threshold $\tau=\Theta(\log|\text{Atom}(S_{O})|)$ grows logarithmically with knowledge base size.
This policy generalizes classical caching (e.g., LFU) by incorporating computational depth $Dd(q|S_{O})$ and semantic information $H_{q}$, and explains the phase transition in storage utility observed empirically.
V-B Phase Transition
Theorem 12 (Storage Utility Phase Transition).
Define $\alpha=E_{\text{storage}}/E_{\text{compute}}$. There exists critical point
$\alpha_{c}=\frac{H(Q)}{\mathbb{E}[\text{Dd}(Q|S_{O})]\cdot\ln(2)}$ (96)
such that: (i) for $\alpha<\alpha_{c}$, $\partial L/\partial M\sim-\exp(-\alpha/\alpha_{c})$; (ii) for $\alpha>\alpha_{c}$, $\partial L/\partial M\sim-1/M$.
@@ VI Numerical Validation and Simulations
All datasets satisfy $H(q\mid S_{O})\gg\log|\mathrm{Atom}(S_{O})|$ (verified empirically), placing them firmly in the information-rich regime. We validate our theoretical predictions through extensive numerical experiments on synthetic knowledge systems. Knowledge bases range from 1K to 100K atoms. Query distributions include uniform and Zipfian ($\alpha\in\{1.0,1.2,1.5,1.8,2.0\}$). Physical parameters are: temperature T=300T=300K.
@@
In Experiment 1, we validate Theorem LABEL:thm:duality (Frequency-Weighted Duality) by testing individual queries sampled from realistic distributions. For each query $q$, we compute its access frequency $f_{q}$ from a Zipfian distribution, then verify that the optimal amortized cost satisfies the theoretical bound $\text{Cost}_{\text{amort}}(q)\geq H_{q}(1+\frac{1}{c\log|K|}-\frac{1}{f_{q}})$. We test 50 randomly sampled queries from each of three knowledge bases: Small (1K atoms), Medium (10K atoms), Large (100K atoms), with average logical depths 3–5.
TABLE II: Single-Query Duality Theorem Validation (Theorem LABEL:thm:frequency_weighted_duality)
KB Size Entities # Queries Sat. Rate Avg. Margin
Small 1.00K 50 100.0% 6,785.7
Medium 10.00K 50 100.0% 280,828.3
Large 100.00K 50 100.0% 10,329,841.7
@@
Note: Sat. Rate = fraction of queries satisfying $\mathrm{Cost}_{\mathrm{amort}}(q)\geq H_{q}\bigl(1+\frac{1}{c\log|K|}-\frac{1}{f_{q}}\bigr)$. Avg. Margin = average difference between observed cost and theoretical bound.
The results demonstrate perfect validation of Theorem 5: all 150 tested queries (50 per knowledge base) satisfy the frequency-weighted duality bound. The large positive margins (6.8K–10.3M nats) indicate that the bound is not tight in practice, which is expected since the theorem provides a fundamental lower limit rather than an achievable optimum. The increasing margin with knowledge base size reflects the $1/f_{q}$ term in the bound—larger knowledge bases have more queries with lower individual frequencies, leading to more conservative bounds.
@@
In Experiment 2, we vary storage capacity from 0
Figure 1: Phase transition in query latency. The figure shows: (top-left) latency vs. storage capacity, (top-right) gradient $\partial L/\partial M$ marking the phase transition signature, (bottom-left) cache hit rate, and (bottom-right) latency vs. storage in bits on log scale. The red dashed vertical lines mark the transition point at approximately 10% storage capacity, where latency improvement shifts from rapid to gradual, demonstrating the phase transition phenomenon predicted by theory.
Results (Figure 1): The phase transition occurs at approximately 10% of the query set, identified by the inflection point in the latency gradient. Below this point, increasing storage yields rapid latency improvement (steep gradient); above it, further storage provides diminishing returns (gradual gradient). The transition validates the theoretical prediction that storage utility undergoes a qualitative shift at a critical capacity determined by information-theoretic and workload parameters.
@@
TABLE III: Baseline Comparison at Cache Size = 50 (5% of queries)
Strategy Hit Rate Latency Improv.
LRU 87.2% 1.54 baseline
LFU 87.9% 1.50 2.6%
TrueMI 51.6% 2.96 -92%
FreqDepth (Ours) 91.5% 1.36 11.6%
Figure 2: Baseline strategy comparison across cache sizes (10-500). The three panels show: (left) cache hit rate, (center) average query latency, and (right) total computation cost. Our FreqDepth strategy (freq $\times$ depth heuristic) consistently outperforms LRU and LFU across all metrics and cache sizes. TrueMI (mutual information maximization) performs poorly due to ignoring access frequency.
At the constrained cache size of 50 (Table 2), our FreqDepth strategy achieves 91.5% hit rate vs. 87.2% for LRU, with 11.6% latency improvement. Figure 2 shows performance across the full range (cache sizes 10-500): FreqDepth consistently outperforms LRU/LFU at all sizes. The theoretically optimal TrueMI performs poorly (51.6% hit rate at cache=50) because it greedily selects queries that maximize global mutual information without considering access frequency—it caches informationally rich but rarely accessed queries. In contrast, FreqDepth’s $\text{freq}\times\text{depth}$ heuristic balances information gain with access probability, achieving practical superiority. At larger cache sizes (e.g., cache $\geq$ 300), all classical strategies converge to $95\%$ hit rates as capacity constraints relax.
In Experiment 4, we test robustness via sensitivity analysis across: (i) five Zipf exponents $\alpha\in\{1.0,1.2,1.5,1.8,2.0\}$ modeling query frequency distributions; (ii) five average logical depths $\{2,3,5,7,10\}$; (iii) four knowledge base scales from 100 to 5000 entities:
Figure 3: Sensitivity analysis. Left pair: cost vs. storage fraction for different Zipf parameters $\alpha$ and optimal $\beta^{*}$ vs. $\alpha$. Right pair: cost vs. storage fraction for different logical depths and optimal $\beta^{*}$ vs. depth. Theoretical bounds hold across all configurations.

### v4

@@ IV-C Entropy Production and Irreversibility
The energy bounds established in this section can be recast in the language of entropy production, providing a complementary perspective rooted directly in the second law of thermodynamics.
IV-C1 Entropy Production in Query Answering
Consider a knowledge system operating in thermal equilibrium at temperature $T$. When the system answers a query $q$ by deriving it from a knowledge base $\mathcal{S}_{O}$, the computation constitutes an irreversible process that produces entropy in the environment.
Proposition 6 (Entropy Production Bound for Query Answering).
Let $\Pi=(S,C,R)$ be a mixed strategy answering a query $q$ from knowledge base $\mathcal{S}_{O}$ in the classical thermodynamic regime (Assumption 3). The minimum entropy produced in the environment satisfies:
$\Delta S_{\mathrm{env}}(q)\geq k_{B}\ln 2\cdot D_{d}(q\mid\mathcal{S}_{O}\cup S),$ (127)
or equivalently, in terms of derivation entropy:
$\Delta S_{\mathrm{env}}(q)\geq k_{B}\cdot H_{\mathrm{derive}}(q\mid\mathcal{S}_{O}\cup S).$ (128)
By Proposition 4, answering query $q$ requires energy $E_{\mathrm{compute}}(q)\geq D_{d}(q\mid\mathcal{S}_{O}\cup S)\cdot k_{B}T\ln 2$. In the quasistatic limit of the classical thermodynamic regime, this energy is dissipated as heat $Q_{\mathrm{diss}}=E_{\mathrm{compute}}$ into the thermal environment. The corresponding entropy increase in the environment is
$\Delta S_{\mathrm{env}}=\frac{Q_{\mathrm{diss}}}{T}\geq\frac{D_{d}(q\mid\mathcal{S}_{O}\cup S)\cdot k_{B}T\ln 2}{T}\\ =D_{d}(q\mid\mathcal{S}_{O}\cup S)\cdot k_{B}\ln 2.$ (129)
Substituting $H_{\mathrm{derive}}(q\mid\mathcal{S}_{O}\cup S)=D_{d}(q\mid\mathcal{S}_{O}\cup S)\cdot\ln 2$ from Corollary 3 yields (128). ∎
Remark 14 (Duality with Proposition 4).
Proposition 6 is the entropic reformulation of Proposition 4: the energy bound $E\geq D_{d}\cdot k_{B}T\ln 2$ and the entropy bound $\Delta S_{\mathrm{env}}\geq D_{d}\cdot k_{B}\ln 2$ are related by the fundamental thermodynamic identity $\Delta S_{\mathrm{env}}=E/T$ in the quasistatic regime. This duality reflects the interchangeability of energy and entropy as thermodynamic costs at fixed temperature.
IV-C2 Expected Entropy Production and Mutual Information
For a query distribution $P_{Q}$, we can relate the expected entropy production to the conditional mutual information $I(S;Q\mid\mathcal{S}_{O})$ introduced in Section III-D.
Proposition 7 (Expected Entropy Production Bound).
Under the conditions of Proposition 6, let $P_{Q}$ be a query distribution satisfying the information-rich regime (Definition 8) and the uniform polynomial-depth regime (Assumption 4). The expected environmental entropy production satisfies:
$\mathbb{E}_{q\sim P_{Q}}\bigl[\Delta S_{\mathrm{env}}(q)\bigr]\\ \geq\frac{c\cdot k_{B}}{L_{S}}\cdot\bigl(H(Q\mid\mathcal{S}_{O})-I(S;Q\mid\mathcal{S}_{O})\bigr)-O(k_{B}),$ (130)
where $c>0$ is the constant from Theorem 4, $L_{S}=\log(|\mathrm{Atom}(\mathcal{S}_{O}\cup S)|+D_{d}^{\sup})$ is the effective complexity parameter, and $I(S;Q\mid\mathcal{S}_{O})$ is the conditional mutual information between storage $S$ and query $Q$ given $\mathcal{S}_{O}$.
By Proposition 6, $\mathbb{E}[\Delta S_{\mathrm{env}}(q)]\geq k_{B}\cdot\mathbb{E}[H_{\mathrm{derive}}(q\mid\mathcal{S}_{O}\cup S)]$. Applying inequality (116) from the proof of Theorem 6 yields the claimed bound. ∎
This result provides a thermodynamic interpretation of storage utility: each nat of mutual information $I(S;Q\mid\mathcal{S}_{O})$ reduces expected entropy production by $c\cdot k_{B}/L_{S}$. Storage that anticipates future queries thus reduces not only computational cost but also environmental entropy production.
IV-C3 Irreversibility and the Limits of Reversible Computation
Remark 15 (Reversible Computation).
The entropy production bounds above apply to logically irreversible computation. Bennett’s reversible simulation theorem [bennett1973logical] shows that any irreversible computation can be made reversible with polynomial overhead in time and logarithmic overhead in space. For knowledge systems, this implies that the Landauer cost could in principle be avoided by storing $O(D_{d}\log D_{d})$ bits of auxiliary information, assuming the original derivation uses time and space proportional to $D_{d}$. However, this shifts cost from computation to storage without eliminating the fundamental information-theoretic constraints of Theorem 5. The practical relevance of reversible computation for knowledge systems remains limited by engineering challenges and the overhead of reversible simulation.
V Design Rules for Hybrid Storage–Computation Strategies
We now move from the instance-wise and information-theoretic analysis of Section III to higher-level design rules for hybrid strategies that combine storage and computation. Rather than introducing new formal theorems, our goal here is to distill the implications of Theorem 5, the mutual-information analysis of Section III-D, and the thermodynamic constraints of Section IV into simple decision guidelines.
V-A Frequency-Dependent Choice of Storage vs. Computation
The frequency-weighted tradeoff in Theorem 5 provides a natural criterion for deciding, for each query type $q$, whether it should be answered purely by on-demand derivation, purely by storage (caching), or by a hybrid strategy that stores partial information.
Recall the notation from Section III-B:
  • $H_{q}:=H(q\mid S_{O})=K(\text{enc}{q}\mid\text{enc}{\text{Atom}(S_{O})})\ln 2$ is the conditional algorithmic entropy of $q$ given $S_{O}$.
  • $f_{q}:=N\cdot P_{Q}(q)$ is the expected number of accesses of query $q$ over a horizon of $N$ queries (Definition 9).
  • For a storage structure $S$, we denote $m_{S}:=|\text{Atom}(S_{O}\cup S)|$ and $n_{S}:=Dd(q\mid S_{O}\cup S)$, and define
    $L_{S}:=\log(m_{S}+n_{S}),$ (131)
    which plays the role of an effective logarithmic compression factor in Theorem 4 and 5.
Under the assumptions of Theorem 9, there exist constants $C,C>0$ such that for any admissible hybrid strategy $\Pi=(S,C,R)$ serving a fixed query type $q$ we have
$\mathrm{Cost}_{\mathrm{amort}}(\Pi;q)\;=\;\frac{|S|}{f_{q}}+H_{\mathrm{derive}}(q\mid S_{O}\cup S)\\ \;\leq\;\min\left\{\frac{H_{q}}{f_{q}},\;C\,\frac{H_{q}}{L_{S}}\right\}+O(1),$ (132)
where all costs are measured in nats (or bits up to a constant factor).
The two competing terms inside the minimum correspond to two extremal—and constructive—strategies:
  1. 1.
    Pure storage (caching) strategy: store a near-optimally compressed representation of the answer to $q$ of length at most $H_{q}+O(1)$ nats (or equivalent bits) so that $H_{\mathrm{derive}}(q\mid S_{O}\cup S)=O(1)$. This yields an amortized per-access cost $\mathrm{Cost}_{\mathrm{amort}}\lesssim H_{q}/f_{q}+O(1)$.
  2. 2.
    Pure computation (no auxiliary storage) strategy: take $S=\varnothing$ and derive $q$ from $S_{O}$ on demand. By Theorem 4, this yields an amortized cost $\mathrm{Cost}_{\mathrm{amort}}\lesssim CH_{q}/L_{S^{(0)}}+O(1)$, where
    $L_{S^{(0)}}=\log\bigl(|\text{Atom}(S_{O})|+Dd(q\mid S_{O})\bigr)$
    is the baseline compression factor for inference from $S_{O}$ alone.
Comparing these two extremes leads to a natural critical frequency scale. In the baseline case $S=\varnothing$ we define
$f_{c}^{(0)}(q)\;\asymp\;L_{S^{(0)}}\;=\;\log\bigl(|\text{Atom}(S_{O})|+Dd(q\mid S_{O})\bigr).$ (133)
When $f_{q}$ is much smaller than $f_{c}^{(0)}(q)$, the term $H_{q}/f_{q}$ is larger than $CH_{q}/L_{S^{(0)}}$, so it is more efficient to answer $q$ by on-demand derivation from $S_{O}$ (pure computation). Conversely, when $f_{q}$ is much larger than $f_{c}^{(0)}(q)$, the caching term $H_{q}/f_{q}$ dominates the pure computation cost, so precomputing and storing (some representation of) the answer to $q$ becomes more attractive.
This motivates the following rule-of-thumb.
Policy 1 (Frequency-aware caching rule (informal)).
Fix a base knowledge $S_{O}$, a query type $q$, and its expected access frequency $f_{q}=N\cdot P_{Q}(q)$ over a horizon of $N$ queries. Let
$f_{c}^{(0)}(q)\;\asymp\;\log\bigl(|\text{Atom}(S_{O})|+Dd(q\mid S_{O})\bigr)$
be the baseline critical frequency as in (133). Then, up to multiplicative constants and lower-order terms:
  • If $f_{q}\ll f_{c}^{(0)}(q)$, the pure computation strategy achieves amortized cost $\Theta(H_{q}/L_{S}^{(0)})$, which is better than the pure storage cost $\Theta(H_{q}/f_{q})$.
  • If $f_{q}\gg f_{c}^{(0)}(q)$, it is typically more efficient to cache the answer to $q$ (or suitable intermediate results), so that most of the cost is paid once in storage and the marginal per-access cost becomes $O(1)$.
  • If $f_{q}\approx f_{c}^{(0)}(q)$, hybrid strategies that store partial information about $q$ (e.g., indices, intermediate derivations) can provide a good balance between storage footprint and computation cost.
Remark 16 (Semantic richness and normalized thresholds).
The entropy $H_{q}$ and the depth $Dd(q\mid S_{O})$ together capture how “semantically rich” and how “derivationally deep” a query type $q$ is with respect to the base knowledge $S_{O}$. A convenient normalized measure is
$\rho(q\mid S_{O})\;:=\;\frac{Dd(q\mid S_{O})}{H_{q}},$
which roughly measures the average derivation depth per nat of semantic content. The ratio $\rho(q\mid\mathcal{S}O)$ measures derivation depth per unit of semantic content. In the information-rich regime, the normalized per-nat derivation cost $H{\text{derive}}/H_{q}$ is governed by $1/L_{S}=1/\log(m+D_{d})$, which decreases slowly as $D_{d}$ grows. However, the absolute derivation cost $D_{d}\cdot k_{B}T\ln 2$ in energy units (Proposition 4) increases with $D_{d}$. Thus, for fixed $H_{q}$, queries with large $\rho$ incur higher absolute energy costs per derivation, making caching more attractive in energy-constrained settings. While Policy 1 is only a rule-of-thumb, this normalized perspective provides additional intuition for how semantic structure interacts with access frequency in storage decisions.
V-B System-Level Storage Allocation via Mutual Information
Section III-D analyzed the system-level cost
$\mathbb{E}[\mathrm{Cost}]\;=\;\frac{|S|}{N}+\mathbb{E}_{q\sim P_{Q}}\bigl[H_{\mathrm{derive}}(q\mid S_{O}\cup S)\bigr]$
through the lens of Shannon entropy and conditional mutual information. Interpreting
and recalling that
$H(Q\mid S_{O},S)\;=\;H(Q\mid S_{O})-I(S;Q\mid S_{O}),$
we see that, under a fixed query distribution $P_{Q}$ and base knowledge $S_{O}$, the storage set $S$ is beneficial precisely to the extent that it increases the conditional mutual information $I(S;Q\mid S_{O})$, i.e., it reduces the residual uncertainty about the (random) query $Q$ beyond what is already known from $S_{O}$.
Treating $S$ as a finite subset of a ground set $U$ of candidate storage items (facts, cached answers, intermediate derivations, etc.), Section III-D showed that the function
$f(S)\;:=\;I(S;Q\mid S_{O})$
is nonnegative, monotone, and submodular in $S$ (Proposition 3). As a consequence, under a budget constraint $|S|\leq M_{\max}$ on the total storage, the problem
$\max_{S\subseteq U:\,|S|\leq M_{\max}}I(S;Q\mid S_{O})$
admits efficient (11/e)(1-1/e)-approximation by the standard greedy algorithm (Corollary 5). Informally, the greedy policy repeatedly chooses the storage item with the largest marginal increase in $I(S;Q\mid S_{O})$ per unit of storage cost, subject to the remaining budget.
This leads to the following high-level prescription for system-wide storage allocation.
Policy 2 (Mutual-information-driven storage selection).
Given a base knowledge $S_{O}$, a query distribution $P_{Q}$, and a storage budget $M_{\max}$ (in bits), define the utility of a candidate storage set $S$ by $f(S)=I(S;Q\mid S_{O})$. Then:
  1. 1.
    Approximate the optimal storage allocation by a greedy procedure that iteratively adds the candidate item with highest marginal gain
    $\frac{I(S\cup\{x\};Q\mid S_{O})-I(S;Q\mid S_{O})}{\Delta|S|}$
    per unit increase $\Delta|S|$ in storage, until the budget $M_{\max}$ is exhausted.
  2. 2.
    The resulting $S_{\mathrm{greedy}}$ is guaranteed to achieve at least a (11/e)(1-1/e) fraction of the optimal mutual-information gain $I(S^{\star};Q\mid S_{O})$, and hence to reduce the expected inference burden $\mathbb{E}[H_{\mathrm{derive}}(q\mid S_{O}\cup S)]$ (and the associated computing energy) nearly as well as any other storage allocation of the same size, up to this universal constant factor.
Policy 2 complements the per-query frequency rule in Policy 1: while the latter identifies, for each $q$, whether caching is worthwhile in principle, the former tells us how to prioritize and share limited storage capacity across many query types so as to maximize global benefit.
V-C Practical Implications and Phase-Transition Intuition
The theoretical results described above, together with Policies 1 and 2, have direct implications for the design of large-scale knowledge systems.
Parametric memory in large language models
For a model with $N$ effective parameters, we can view its weights as encoding a storage set $S$ with capacity on the order of $N$ effective bits. Facts or patterns that are both (i) derivationally deep with respect to the model’s base world-knowledge $S_{O}$, and (ii) frequently queried (high $f_{q}$), are natural candidates to be stored in parameters, whereas shallow or rare facts are better left to on-demand retrieval or tool use. Policies 1 and 2 suggest that an optimal allocation of parametric capacity resembles a mutual-information-maximizing cache adapted to the distribution of downstream tasks.
Materialized views in knowledge graphs
In graph databases, multi-hop join patterns $q$ with large $Dd(q\mid S_{O})$ but modest algorithmic entropy $H_{q}$ (e.g., repeated schema-driven patterns) merit precomputation into materialized views when their access frequencies exceed the logarithmic critical scale $f_{c}^{(0)}(q)$. The submodularity of $I(S;Q\mid S_{O})$ implies that precomputing overlapping views exhibits diminishing returns, and a greedy selection of views by marginal contribution to $I(S;Q\mid S_{O})$ is near-optimal.
Edge caching under energy constraints
On edge devices, storage is energetically costly to maintain (Proposition 5), while remote computation or cloud queries incur network and latency costs. Balancing $E_{\mathrm{storage}}$ against $E_{\mathrm{compute}}$ in this context amounts to choosing a storage budget $M_{\max}$ and a set $S$ such that the expected total energy per query (including both maintenance and computation) is minimized. Our framework shows that, in the information-rich regime, there is a qualitative transition in the optimal policy as the relative cost of storage versus computation varies: when storage is cheap, it is optimal to cache many $f_{q}, high$-DdDd queries; when storage is expensive, it is preferable to serve most queries by on-demand computation or remote access.
We deliberately refrain from formalizing this last observation as a sharp “phase-transition theorem”, since its precise shape depends on hardware- and workload-specific constants (e.g., energy per bit of storage, energy per inference step, query arrival statistics). The qualitative picture, however, is robust: as the energetic cost of storage increases relative to computation, the optimal allocation of storage mass shifts from aggressively caching frequent queries to favoring on-demand computation, with a crossover regime governed by the logarithmic critical frequencies (133) and by the mutual-information structure of $(S,Q\mid S_{O})$.
@@ VI Numerical Validation and Simulations
To ensure our experiments satisfy the information-rich regime assumption (Definition 8: $H(q\mid S_{O})\gg\log|\mathrm{Atom}(S_{O})|$), we fix the knowledge base size at 100K atoms across all main experiments and vary only the query set size. Empirical validation confirms that the information-rich regime begins at approximately $n\geq 50{,}000$ atoms. Query distributions follow Zipfian ($\alpha=1.5$). Physical parameters: temperature T=300T=300K.
In Experiment 1, we validate Theorem 5 (Frequency-Weighted Storage-Computation Tradeoff) through two complementary analyses:
(1a) N-Sweep Analysis: We vary the query horizon $N$ from 100 to 50,000 to verify the theoretical prediction that as $N$ increases: (i) storage amortized cost $H_{q}/f_{q}=H_{q}/(N\cdot p_{q})$ decreases, (ii) more queries become worth storing (optimal $\beta^{*}$ increases), and (iii) total optimal cost decreases. Figure 1 shows the results on a 100K-atom knowledge base with 1000 queries:
Figure 1: N-sweep analysis validating Theorem 5. Top-left: Amortized costs (storage, computation, optimal) vs NN. Top-right: Optimal storage fraction $\beta^{*}$ increases from 0% to 92% as $N$ grows. Bottom-left: Bound tightness remains stable. Bottom-right: Number of queries worth storing increases with $N$.
Key findings: Optimal $\beta^{*}$ ranges from 0% ($N=100$) to 91.8% ($N=50,000$), confirming that larger query horizons favor more aggressive caching. The optimal cost decreases from 4.78 to 0.69 (86% reduction), validating the theoretical bound’s prediction that amortization improves with $N$.
@@
(1b) Fixed-N Validation: For each query $q$, we compute its access frequency $f_{q}=N\cdot P_{Q}(q)$ from a Zipfian distribution ($\alpha=1.5$), then verify that the optimal per-query cost satisfies the theoretical upper bound $\text{Cost}_{\text{amort}}(\Pi;q)\leq\min\{H_{q}/f_{q},C\cdot H_{q}/L_{S}\}+O(1)$. We test three query set sizes on the same 100K-atom knowledge base: Small (100 queries), Medium (1000 queries), Large (10000 queries), with average derivation depths 3–5.
TABLE II: Theorem 5 Validation: Upper Bound Satisfaction
KB Size Entities # Queries Sat. Rate Avg. Tightness High-freq Tightness
Small 100K 100 100.0% 0.81 0.89
Medium 100K 1000 100.0% 0.53 0.89
Large 100K 10000 100.0% 0.50 0.89
@@
Note: Sat. Rate = fraction of queries satisfying $\mathrm{Cost}_{\mathrm{amort}}(\Pi;q)\leq\min\{H_{q}/f_{q},C\cdot H_{q}/L_{S}\}$ (with $C=2$). Tightness = actual cost / theoretical bound (closer to 1.0 = tighter bound).
The results demonstrate perfect validation of Theorem 5: all tested queries across three knowledge bases satisfy the frequency-weighted upper bound. The average tightness of 0.50–0.93 indicates that the bound is reasonably tight, especially for high-frequency queries (tightness $\approx 0.89$). The lower tightness for low-frequency queries reflects the constant factor $C=2$ in the bound; with $C=1$, the tightness would approach 1.0. This validates that the theoretical bound correctly captures the storage-computation tradeoff with appropriate constant factors.
We also validate Theorem 6 (Information Lower Bound) by computing the minimum achievable cost $\bar{I}(\Pi)$ across all storage fractions and comparing to the theoretical lower bound $c\cdot H(Q|S_{O})/L_{S}$. Using $c=0.1$ as a conservative estimate:
TABLE III: Theorem 6 Validation: Lower Bound Analysis
KB Queries $\min\bar{I}$ Bound Satisfied Ratio
[…]
(1c) Deep Analysis of Core Definitions: We validate three foundational elements of our theoretical framework by varying $m=|\mathrm{Atom}(S_{O})|$ from 100 to 100,000:
  • Theorem 4 (Derivation Entropy Scaling): The theorem states $H_{\mathrm{derive}}=\Theta(H_{q}/L_{S})$. We compute the ratio $H_{\mathrm{derive}}\cdot L_{S}/H_{q}$ for each query across different $m$ values. Results show this ratio is remarkably stable: mean =0.217=0.217, range $[0.198,0.228]$. This confirms the $\Theta(1)$ constant factor is approximately 0.22.
  • Derivation Depth Distribution: The polynomial-depth assumption (Assumption 4) requires $D_{d}\leq\mathrm{poly}(m)$. Our experiments show $D_{d}^{\max}$ ranges from 3 to 7 as $m$ increases from 100 to 100,000, with $D_{d}/\log(m)\approx 0.6$ remaining stable. This validates that derivation depth scales logarithmically with knowledge base size for our query classes.
  • Effective Complexity Factor: The parameter $L_{S}=\log(m+n)$ closely tracks $\log(m)$ across all configurations, with $L_{S}/\log(m)\in[1.0,1.1]$, confirming the polynomial-depth regime where $n=D_{d}=O(\log m)$.
Figure 2: Deep analysis of core theorems. Top row: (left) Theorem 1 ratio $H_{\mathrm{derive}}\cdot L_{S}/H_{q}$ vs mm—stable around 0.22; (center) $H_{\mathrm{derive}}$ vs $H_{q}/L_{S}$—linear relationship; (right) $L_{S}$ vs $\log(m)$. Bottom row: (left) $D_{d}$ range vs $m$; (center) $D_{d}^{\max}/\log(m)$ stability; (right) summary metrics.
@@
In Experiment 2, we vary storage capacity from 0% to 100% of the query set and measure actual query cost to validate the predicted phase transition. We test on the 100K-atom knowledge base with varying query set sizes and compute the optimal storage fraction via marginal cost-benefit analysis.
Figure 3: Phase transition analysis for Medium query set (100K atoms, 1000 queries). Top-left: Cost vs. storage fraction with theoretical and empirical transition points marked. Top-right: Cost gradient showing phase transition signature. Bottom-left: Query frequency $f_{q}$ vs. critical frequency $f_{c}$ identifying the transition. Bottom-right: Frequency regime distribution (high-freq/critical/low-freq). The theoretical transition point is determined by marginal cost-benefit analysis.
Results (Figure 3): The theoretical optimal storage fraction $\beta^{*}$ is computed via marginal cost-benefit analysis: store query $q$ if the marginal benefit $p_{q}\cdot D_{d}(q)\cdot\lambda$ exceeds the marginal cost $1/N$. For the Medium query set (1000 queries), this yields $\beta^{*}=72.3\%$ (723 queries worth storing), matching the empirical optimum of $75.5\%$ with only $3.2\%$ difference. For the Large query set (10000 queries), theoretical $\beta^{*}=7.1\%$ vs. empirical $8.2\%$ (1.0% difference). The frequency regime analysis reveals 30 high-frequency queries ($f_{q}>2f_{c}$), 47 critical queries ($f_{q}\approx f_{c}$), and 923 low-frequency queries ($f_{q}<f_{c}/2$). This validates that the marginal analysis correctly predicts the optimal storage allocation.
@@
TABLE IV: Baseline Comparison at Cache Size = 50 (5% of queries)
Strategy Hit Rate Cost Improv.
LRU 87.2% 0.50 baseline
LFU 87.9% 0.52 -4%
TrueMI 51.3% 1.98 -296%
FreqDepth (Ours) 91.5% 0.37 26%
Figure 4: Baseline strategy comparison across cache sizes (10-200). Left: Average cost (paper-defined: $|S|/N+\mathbb{E}[H_{\mathrm{derive}}]$) vs cache size. Right: Cache hit rate vs cache size. Our FreqDepth strategy (freq $\times$ depth heuristic) consistently outperforms LRU and LFU across both metrics. TrueMI (mutual information maximization) performs poorly due to ignoring access frequency.
At the constrained cache size of 50 (Table IV), our FreqDepth strategy achieves 91.5% hit rate vs. 87.2% for LRU, with 26% cost reduction. Figure 4 shows performance across the full range (cache sizes 10-200): FreqDepth consistently outperforms LRU/LFU at all sizes. The TrueMI strategy performs poorly (51.3% hit rate at cache=50) because it greedily selects queries that maximize global mutual information without considering access frequency—it caches informationally rich but rarely accessed queries. This validates our theoretical insight that mutual information maximization $\neq$ cost minimization. In contrast, FreqDepth’s $\text{freq}\times\text{depth}$ heuristic balances information gain with access probability, achieving practical superiority.
In Experiment 4, we test robustness via sensitivity analysis across: (i) five Zipf exponents $\alpha\in\{0.5,1.0,1.5,2.0,2.5\}$ modeling query frequency distributions; (ii) seven average derivation depths $\{1,2,3,5,7,10,15\}$; (iii) knowledge base scales from 50 to 5000 entities:
Figure 5: Sensitivity analysis. Top row (zipf_regimes_v4.pdf): Left panel shows frequency regime distribution (high-freq, critical, low-freq) vs Zipf parameter $\alpha$; right panel shows optimal storage fraction $\beta^{*}$ vs $\alpha$. Bottom row (sensitivity_depth_v4.pdf): Left panel shows cost vs storage fraction for different derivation depths; right panel shows optimal $\beta^{*}$ vs depth. Theoretical bounds hold across all configurations.
Results show broad applicability: as Zipf parameter $\alpha$ increases from 0.5 to 2.5, the fraction of high-frequency queries decreases from 100% to 6%, and the optimal storage fraction $\beta^{*}$ decreases from 100% to 30%. This validates the theoretical prediction that query skewness determines the optimal storage strategy—more skewed distributions (larger $\alpha$) benefit less from full caching. Scale tests confirm $H(Q)\sim\log(n)$ scaling: fitting yields $H(Q)\approx 0.42\log(n)+0.62$ bits across the 50–5000 entity range, with the ratio $H(Q)/\log(n)$ remaining stable around 0.53.
Figure 6: Entropy scaling validation. Left: $H(Q)$ vs. knowledge base size $n$ on log scale, with fitted line $H(Q)\approx 0.42\log(n)+0.62$. Right: Ratio $H(Q)/\log(n)$ remains stable around 0.53, confirming the theoretical $H(Q)\sim\log(n)$ scaling.
Summary of Experimental Validation.
Table V summarizes the validation status of our main theoretical results:
TABLE V: Summary of Theoretical Validation
Theorem Status Key Finding
Thm 5 (Upper Bound) 100% satisfaction